OpenStudy (sette):

A company builds computers. It costs $7,600 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? c(x)= 460x-3000 c(x)= 3000x-460 c(x)= 460x+3000 c(x)= 3000x+$460

OpenStudy (bahrom7893):

plug in y = 7,600 and x = 10 and y = 12,200 and x = 20 to check

OpenStudy (sette):

Okay !

OpenStudy (sette):

wait to all of them ?

OpenStudy (anonymous):

7600=10*a+b 12200=20*a+b find a and b

OpenStudy (anonymous):

y=460x+3000

OpenStudy (anonymous):

c(x)= 460x+3000